Find the distinct interval of length 1 containing a root or solution of f(x)= x ³-3x+5 using IVT
1
Expert's answer
2021-10-20T07:00:04-0400

"f(x)= x^3-3x+5
\\f(-2)=(-2)^3-3\times(-2)+5=-8+6+5=3>0
\\f(-3)=(-3)^3-3\times(-3)+5=-27+9+5=-13<0"

As "f(-2)>0 , f(-3)<0"

Then from IVT, there will be a root between "[-2,-3]."

This is only interval of length 1 for which , we will get the roots or solution of "f(x)."